OpenStudy (anonymous):

please help!!! important!!!!!!!!!! Answer is A Is x - 8 a factor of the function f(x) = -2x3 + 17x2 - 64? Explain. (2 points) Answer Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64. No. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is not a factor of f(x) = -2x3 + 17x2 - 64. Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is not zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 6

OpenStudy (anonymous):

Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64.

OpenStudy (anonymous):

here f(8)=-1024 +1088 -64=0

OpenStudy (anonymous):

the answer is A ???

OpenStudy (anonymous):

you can sub the x-value that makes x-8 equal zero, into the function. so the x-value that makes x-8=0 is 8, correct? so if you agree, then we can sub in 8 into the function. if you sub in 8 into the function and you get a value of zero, then x-8 is a factor of the function.